OpenStudy (twalt13):

please can anyone help me with my question !!! Use basic identities to simplify the expression. sinx cosx secx cscx

OpenStudy (nubeer):

\[secx = \frac{ 1 }{ cosx }\] \[cosecx = \frac{ 1 }{ sinx}\] now just plug these in your question \[sinx \times cosx \times \frac{ 1 }{ cosx } \times \frac{ 1 }{ sinx }\]

OpenStudy (twalt13):

is it 1 ?

OpenStudy (nubeer):

yes

OpenStudy (twalt13):

thank you